OpenStudy (anonymous):

X^1-X^-1+x=2 A.0 B.1/2 C.11/2 D.1/2

OpenStudy (raden):

can u draw your equation above, please

OpenStudy (anonymous):

idk how to do this.. i need help so bad.. =(

OpenStudy (anonymous):

In the reply window, there is a blue button for equation and one for drawing. Its on the left bottom box where you type your post.

OpenStudy (anonymous):

\[x^1-x^{-1}+x=1\\ 2x-x^{-1}=1\\ 2x-\frac{1}{x}-1=0\\ x\ne 0\\ \mbox{both sides}\times x\\ 2x^2-1-x=0\\ x^2-\frac{x}{2}-\frac{1}{2}=0 \] Now you use the formula to solve for the second degree equation.

OpenStudy (anonymous):

But I just realized. Its easier to just test out the values in the answers: x=0, \(0^1-0^{-1}+0\ne1\) so its not A.

OpenStudy (anonymous):

ahhh okays thanks

OpenStudy (jhannybean):

\[\large 2x^2 -x-1\]\[\large x^2-x-2\]\[\large (x-2)(x+1)\]\[\large (x-\frac{2}{2})(x+\frac12)\]\[\large (x-1)(2x+1)=0\] Hmmm.

OpenStudy (jhannybean):

First off, you have 2 of the same answers.Are you sure these are the right answers?

OpenStudy (anonymous):

x=1/2, \(\displaystyle\left(\frac{1}{2}\right)^1-\left(\frac{1}{2}\right)^{-1}+\frac{1}{2}\) \[\frac{1}{2}-2+\frac{1}{2}=1-2\ne2\] so its not B either.

OpenStudy (anonymous):

Yeah, those repeated options must be wrong.

OpenStudy (jhannybean):

11/2 doesn't equal 2 either.

OpenStudy (jhannybean):

The answers would have been 1 and -1/2.

OpenStudy (anonymous):

Yes. Those are the answers for the second degree equation too. Just like you factored.

OpenStudy (jhannybean):

Mmhmm.

OpenStudy (anonymous):

I think the kid went away to school. He looked like he was doing his homework in a rush.

OpenStudy (jhannybean):

lol -_-

OpenStudy (anonymous):

I think I made a mistake here \[x^1-x^{-1}+x=1\\ 2x-x^{-1}=1\\ 2x-\frac{1}{x}-1=0\\ x\ne 0\\ \mbox{both sides}\times x\\ 2x^2-1-x=0\\ x^2-\frac{x}{2}-\frac{1}{2}=0\]

OpenStudy (jhannybean):

completing the square?

OpenStudy (jhannybean):

ok,you ended up with the same equation though xD

OpenStudy (anonymous):

It should be: \[x^1-x^{-1}+x=2\\ 2x-x^{-1}=2\\ 2x-\frac{1}{x}-2=0\\ x\ne 0\\ \mbox{both sides}\times x\\ 2x^2-1-2x=0\\ x^2-x-\frac{1}{2}=0\]

OpenStudy (anonymous):

I think its still wrong. Should have used pen and paper.

OpenStudy (anonymous):

Oh ok, it looks like it works now.

OpenStudy (jhannybean):

\[\large 2x -\frac1x -2=0\]\[\large x(2x-\frac1x-2=0)\]\[\large 2x^2 -1-2x =0\]\[\large 2x^2 - 2x - 1 = 0\] 0 does not equal 1 when x = 1/2 , -3/ 2 does not equal 0 when x= 11/2 , 97/2 does not = 0.

OpenStudy (anonymous):

Yes, there is no right answer among those provided. Your alternatives are right: either 1 or -1/2.

OpenStudy (jhannybean):

Mmhmm.